Please help with this problem.
Tawny found $1.95 in quarters and dimes in the chair cushions. There were four times as many dimes as quarters. How many dimes and quarters did she find?
This is what I have so far:
0.20d+0.75d+100=1.95
0.95*4=3.8
This is where i got lost.
Thanks for your help,
Janice

Answer by Theo(13342)

let d = number of dimes
let q = number of quarters
.25 * q is the value of q number of quarters
.10 * d is the value of d number of dimes.
.25 * q + .10 * d = 1.95
d = 4 * q
.25 * q + .10 * 4 * q = 1.95
.25 * q + .4 * q = 1.95
.65 * q = 1.95
q = 3
d = 4 * q = 4 * 3 = 12
.25 * q = .25 * 3 = .75
.10 * d = .10 * 12 = 1.20
1.20 + .75 = 1.95
number of quarters is 3 and number of dimes is 12
key to solving the problem is recognizing that d = 4*q and then substituting for d to make the unknown variables q only.
solve for q and then solve for d after you have the value for q.